A projectile is launched with 30m/s at an angle of 60°with the horizontal.(a) find the position of the projectile, and the magnitude and direction of its velocity at t=2s (b) find the time when the ball reaches the highest point.(c) find the horizontal range

1
Expert's answer
2022-05-09T14:49:41-0400

"l=\frac{v^2\sin2\alpha}g=79.5~m,"

"h=\frac{v^2\sin^2\alpha}{2g}=34.4~m,"

"t=\frac{2v\sin\alpha}g=5.3~s."